# PromptRPA: Generating Robotic Process Automation on Smartphones from Textual Prompts

### v1

@@ 3.1. RPA Task Model
Advances in natural language processing and human-centric design have heralded a more dynamic, and context-aware era in RPA, guided by textual prompts. These textual inputs are not only easily formulated but also offer valuable information to the RPA system, such as the task’s goal, steps, related parameters, and specific applications.
Figure 2. Components of an RPA task model.
To integrate text and the execution requirements of RPA, we propose a formalized RPA task model, illustrated in Figure 2. The model comprises the following components:
@@
  • Function Description ($F$): A segment of text in natural language that represents the overarching objective the RPA aims to achieve.
@@
  • Instruction List ($i_{1:n}$): It consists of formalized instructions, derived from natural language text, encoded in a machine-friendly standard language to direct the operation sequence.
@@
  • Operation Sequence ($o_{1:n}$): An ordered set of actions that the RPA performs, often dynamically adjusted based on the textual prompts and the real-time operating environment.
3.2. The Generation of an RPA Task Model
Our primary objective is to generate an RPA task model from a textual prompt $P$, which involves constructing the model’s components by incorporating related knowledge. Figure 3 illustrates the foundational pipeline for model generation, segmented into the following three stages.
Figure 3. The generation process of an RPA task model.
Information Collection: From the textual prompt $P$, we extract useful information and underlying intentions. Incorporating external knowledge including contextual parameters and tutorials, we prepare all relevant natural language information: the model’s overarching function description $F$ and its granular step description $s_{1:n}$.
Instruction Generation: Employing the collected information, we strategize the task’s progression. Given the instruction definition, we convert the step description into a coherent list of standardized instructions $i_{1:n}$.
Operation Mapping: Following the planned instructions and understanding the context of the actual operating environment, we finalize the operations, resulting in the generation of the operation sequence $o_{1:n}$.
3.3. Existing Challenges
There has been prior work on generating executable operation sequences from text, covering parts or even the entire pipeline described above. However, to our knowledge, the emphasis of existing solutions has largely been on validating the feasibility of such conversions. There is room for improvement in their practical effectiveness as observed in real-world application.
@@ 3.3. Existing Challenges
Firstly, real-world application faces the challenge of input and output diversity. Users’ expressions vary greatly, influenced by their varying levels of knowledge. Developers or skilled users can provide clear, standard instructions for specific operations, while novices or elderly users often struggle to articulate even basic functional requirements, sometimes only having a vague concept of the desired outcome (Jin et al., 2022). Most current research (Pasupat et al., 2018; Li et al., 2020a; Pan et al., 2022) supports explicit, single-step commands, such as “click the confirm button” or “input Hello in the edit box”, but falls short in handling more complex tasks like “activate developer mode”, and are particularly challenged by ambiguous requests like “change the appearance”. As for outputs, the operation sequence is not fixed. As shown in Figure 3, the sequence depends not only on the textual guidance in the RPA task model but also on the operating environment. Factors like smartphone model, app version, and page state can all influence the operation sequence. Taking Figure 4 for an example, when a user wants to post a moment, varying page states may require different operations.
@@
Secondly, to our knowledge, the current machine learning models are not always satisfying in terms of semantic recognition of GUI elements, both in accuracy and efficiency. Despite advances in image recognition algorithms, the small granularity and rich styles of GUI icons (Dixon et al., 2014), coupled with their complex semantics (Passini et al., 2008), pose significant challenges. For instance, subtle distinctions can carry different meanings, like distinguishing between a checked and an unchecked checkbox or discerning a progress bar from a scrollbar, which may appear similar. Moreover, the range of semantic concepts that pre-trained classification models can recognize is limited (Liu et al., 2018). Particularly on mobile pages with insufficient accessibility support, where recognizable text is sparse and reliance on image interpretation is high, semantic understanding of the page degrades (Xie et al., 2020; White et al., 2019). Even state-of-the-art algorithms like Pix2Struct (Lee et al., 2022), “fail to capture the meaning of a more complex screen” when only pixel information is accessible. Moreover, semantic analysis of a small icon within a rich context typically requires several seconds (Wang et al., 2023), a duration that is impractical for efficient operation.
Furthermore, once users issue initial commands, they often find themselves excluded from the subsequent execution process. In cases of task failure, as the entire process is a “black box” to them, it remains unclear which part failed or whether the users could have assisted (Castelvecchi, 2016; Guidotti et al., 2018). Beyond failed tasks, human intervention is often crucial for tasks requiring personalized input, such as selecting a flight or choosing a playlist song. Requiring users to provide all the details in initial commands can be cognitively overwhelming. Additionally, the inherent ambiguity in natural language (Winograd, 1972) requires a mutual confirmation process between the AI and the user, with users needing to communicate their intentions clearly (Ni et al., 2020), and the AI having to reveal enough working details to build user trust (Amershi et al., 2019), possibly necessitating multiple interactions for clarification.
Lastly, existing work is not immediately ready for use out-of-the-box. In practical use, the number of applications and their functions desired by users reaches into the billions (Islam et al., 2010). Prior RPA systems developed using closed tutorial collections, such as UGIF (Venkatesh et al., 2022), PixelHelp (Li et al., 2020a), and AITW (Rawles et al., 2023), support only a fraction of these functions. They become inadequate when faced with unsupported functions, mismatched personalized parameters, or GUI updates. In comparison, online tutorials, while not matching the quality of carefully collected closed datasets, far exceed them in terms of quantity. Broadening the scope of tutorial knowledge to include these online resources could better fulfill the diverse functional demands of users.
3.4. Design Guideline
Building on the RPA generation process and the existing challenges we analyzed detailed above, the system design is founded on several key principles aimed at effectively transforming textual prompts into executable operations:
Foster Versatility: In the RPA generation process, versatility entails accommodating a wide range of inputs and operating environments (outputs). For inputs, the system should address uncertainties, incompleteness, form diversity, and the broad scope of user needs. During information collection, the system should supplement missing information in user expressions, either through predictive analysis or by querying users. For outputs, challenges from the operating environment like rapidly updated pages should be addressed. Hence, during operation mapping, the system requires a deeper understanding of the real-time environment and the ability to make adaptive decisions, not limited to pre-defined instructions.
Optimize Knowledge Utilization: Each stage demands its specific expertise, which should be as exhaustive and open-ended as possible, rather than confined to a predefined set. During information collection, the system should recognize the function needs and retrieve relevant external tutorials if necessary. During instruction generation, adding rules to convert text into instructions enhances the extraction of instructive content and minimizes information loss. During operation mapping, supplementing with real-time page semantics and app-specific information is crucial.
Incorporating User Involvement: While user oversight can identify issues and help enhance system effectiveness, it is crucial to streamline interactions to minimize user burden. A stage-wise generation process helps make the results of each stage transparent and allows for user intervention with minimal effort when necessary, such as through quick quality assessments of text, simple text modifications, or demonstrating a single step with guidance.
Enable Iterative Improvement: System enhancement stems from both capability improvement and knowledge accumulation. Capability improvement relies on mechanisms and models that support iterative optimization. Knowledge accumulation arises from rising usage or manually introduced knowledge supplements. Additionally, addressing knowledge obsolescence is essential, a common issue in outdated mobile tutorials.
4. PromptRPA
Realizing the above design guidelines necessitates a nuanced approach to address the multifaceted challenges during each stage of the RPA generation. Therefore, the utility of a multi-agent system framework here becomes evident. By assigning sub-tasks to specialized agents, the framework enhances the system’s capability to handle complex tasks, ensuring a more comprehensive and efficient task execution process.
To optimize the performance of each agent, we emphasize two core factors: the comprehensiveness of the knowledge base and the stability of decision-making. They provide each agent with a shared pool of knowledge and clear behavioral guidelines, ensuring that they collaborate effectively to accomplish tasks. In alignment with this, Figure 5 illustrates how the system is partitioned into specialized agents and a unified knowledge base. We will delve into each in the following subsections.
@@ 4. PromptRPA
Figure 5. System design of PromptRPA. The blue solid lines indicate the main flow of the RPA generation process, the grey dashed lines represent optional user intervention methods, the purple dashed lines signify knowledge accumulation after successful task execution, and the red solid line denotes instances where the agent requests user intervention.
@@ 4.1. Multi-Agent Design
Our system design is based on the prior agent system framework (Panait and Luke, 2005; Dorri et al., 2018). Drawing inspiration from human capabilities, each agent within our system is endowed with the following components:
  • Memory: The place where past experiences and learned knowledge are stored.
  • Tools: Instruments or utilities the agent utilizes to process data or perform tasks.
  • Strategy: The agent’s internal cognitive mechanism for decision-making.
  • Action: The tasks the agent performs based on its cognition.
Besides these intrinsic components, transparency and user oversight are cornerstones of PromptRPA. The agents are integrated seamlessly with a user interface, allowing users to observe, modify, or augment agent decisions. The interface is not just for monitoring purposes but aims to foster human-AI collaboration.
According to the generation process of an RPA task model detailed in Section 3.2, we have designed six distinct agents specifically tailored to address these three stages.
  • Information Collection:
    • Analysis Agent: Extracts and analyzes information encapsulated in the textual prompt to construct a complete function description $F$.
    • Retrieval Agent: Acquires external knowledge relevant to the function description and merges it with the textual prompt to build an enriched step description $s_{1:n}$.
  • Instruction Generation :
    • Parsing Agent: Transforms the collected information into a series of instructions $i_{1:n}$.
  • Operation Mapping:
    • Grounding Agent: Predicts and eventually executes operations $o_{1:n}$ on the smartphone based on the generated instructions.
    • Mobile Semantics Agent: Extracts valuable semantic information from the mobile interface to aid in operation mapping.
    • Assessment Agent: To balance prediction accuracy and user effort, it reviews the predicted operations, evaluating if human intervention is required.
@@ 4.1.1. Analysis Agent
The agent aims to extract task-related information from the textual prompts, as outlined in Section 3.1. The prompts may include descriptions of the function or specific steps in the model. The agent employs the LLM to segregate the text into these two components.
The function descriptions in the prompts often vary in formality and may lack essential context, such as specific application names or phone models. Figure 6 demonstrates how the agent interacts with users to refine these descriptions. Utilizing the LLM, the agent successively adopts three key strategies:
@@
  • Normalization: The agent rephrases the original function description for clarity and precision.
  • Automated Context Enrichment: The agent utilizes the LLM to determine the essential parameters needed for the RPA task. When parameters are absent in the prompt, the agent consults the context library (detailed in Section 4.2.2) to fetch relevant records as potential candidates.
  • Interactive Context Enrichment: For parameters lacking records in the context library, the agent queries the user to obtain the missing information.
After refining and enriching the function description, the user confirms it via interactive dialogue. The agent then bundles the function description, along with step descriptions (if available), and forwards them to the Retrieval Agent for subsequent processing.
@@ 4.1.2. Retrieval Agent
The Analysis Agent can ensure a clear function description, but it may fall short when it comes to providing detailed step descriptions, which the prompt might not always cover. Addressing this void, the Retrieval Agent is responsible for crafting a complete step description. For efficiency, the agent initially adopts the following strategies:
@@
  • Otherwise, the agent searches the historical RPA repository (detailed in Section 4.2.1) for matching historical RPA task models, using function description as the keyword, with LaBSE (Feng et al., 2022) for comparison, which is a state-of-art multilingual sentence embedding model. If a match is found, the agent will use the matched RPA task model, which has been previously used within our system and is deemed reliable.
If the user is dissatisfied or no match is found, the agent searches for relevant tutorials as the step description by using search engines, with the function description as a query. The tutorial texts from the fetched web pages are refined using SimpRead [^1] https://simpread.pro/’s rule library and summarized by the LLM. Tutorials are then empirically ranked based on criteria including their original search ranking, data source, and textual quality assessed by the LLM. Notably, the alignment with the step description extracted from the user’s prompt is also a factor to consider.
The system presents the top 5 highest-ranking tutorials to users, automatically selecting the highest-ranked tutorial in cases where users do not make an active choice. By this stage, both the function and step description for the RPA are prepared in place.

### v2

@@ 3. Problem of UI Task Automation
UI task automation is traditionally performed by mimicking predefined, rule-based operations on graphical user interfaces (Li et al., 2017). Recent advances in natural language processing and human-centric design have indicated a more dynamic, context-aware era in the task automation, guided by textual prompts. These textual inputs are not only easily formulated but also can offer valuable information, such as the task’s goal, steps, related parameters, and specific applications.
In this section, we introduce Prompt2Task model to better understand the UI task automation process driven by text. Based on this model, we analyze the challenges within the task automation, which inform our system’s architecture.
3.1. Prompt2Task Model
Figure 2. Components of a Prompt2Task model.
Prior task models were developed to provide formalized descriptions of tasks, typically in the context of user interface design. They are primarily used to describe complex system interactions and workflows, such as those involving concurrency and collaboration in high-complexity tasks (Limbourg and Vanderdonckt, 2004). To address task automation on mobile GUI, we propose a simplified UI task model based on ConcurTaskTrees (Paternò et al., 1997), as shown in Figure 2. This model consists of the following core components:
@@ 3.1. Prompt2Task Model
  • Function Description ($F$): A piece of natural language text that represents the overarching objective the automation task aims to achieve.
@@
  • Instruction List ($i_{1:n}$): A list of formalized instructions, derived from natural language text, encoded in a machine-friendly standard language to direct the operation sequence.
@@
  • Operation Sequence ($o_{1:n}$): An ordered set of operations required by the automation task, dynamically adjusted based on the textual prompts and the real-time operating environment.
In the Prompt2Task model, $F$ and $s_{1:n}$ are used to express the task’s goal and steps through natural language. While natural language lacks the rigor of prior task models such as the tree structure of ConcurTaskTrees (Paternò et al., 1997) or the precise descriptions of concurrency, communication, and temporal order in LOTOS (Bolognesi and Brinksma, 1987), it provides sufficient flexibility to describe the core requirements of a task effectively. More importantly, it reduces the complexity for users, making it more accessible for non-experts. Additionally, $i_{1:n}$ and $o_{1:n}$ in the Prompt2Task model are represented as linear lists, since the execution of each automation task is typically carried out on a single device through simulated operations, making task execution inherently linear. Even when task planning involves branching structures, the actual execution remains linear. It does not require the complex structure in Hierarchical Task Analysis (Stanton, 2006) or concurrency mechanisms in Event-Driven Process Chains (Scheer et al., 2005). By generating the operation sequence dynamically, the model is flexible enough to adapt to real-time changes. Therefore, the Prompt2Task model is well-suited to handle the dynamic and diverse requirements of UI task automation efficiently.
3.2. Challenges for UI Task Automation
In this paper, we focus on the process of inputting a textual prompt $P$ to generate the aforementioned Prompt2Task model and ultimately execute the automation task based on the model. The information contained in $P$ is insufficient to accurately generate the model components. Therefore, the process necessitates the integration of additional knowledge, such as application-specific details, user context, and current GUI state. In this process, we identify the following four challenges:
@@ 3.2. Challenges for UI Task Automation
1. Input: The input can be any text related to the task, ranging from formal manuals and tutorials to diverse user inputs. Users’ expressions vary greatly, influenced by their differing levels of knowledge. Developers or skilled users can provide clear, standard instructions for specific operations, while novices or elderly users often struggle to articulate even basic functional requirements, sometimes only having a vague concept of the desired outcome (Jin et al., 2022). Even skilled users might expect to automate tasks with minimal and simple inputs. Most current research (Pasupat et al., 2018; Li et al., 2020a; Pan et al., 2022) accepts explicit, step-by-step commands, such as “click the confirm button” or “input Hello in the edit box”, but falls short in handling more complex commands like “activate developer mode”, and are particularly challenged by ambiguous requests like “change the appearance”.
@@
Table 1. Comparison to prior work. The comparison dimensions include the type of input content, whether the range of supported tasks is limited, whether the operation sequence can be updated in real-time based on the actual UI state, how to handle errors, and what types of knowledge are incorporated.
Prior work Input Unlimited Update Error Knowledge
Tasks Handling
UIBert (Bai et al., 2021) Single step $\times$ ED&R Prepared dataset
Spotlight (Li and Li, 2023) Single step $\times$ ED&R Prepared dataset
Ferret-UI (You et al., 2024; Li et al., 2024b) Single step $\times$ ED&R Prepared dataset
MUG(Li et al., 2024a) Single step $\times$ UI Prepared dataset,
interaction history
AdbGPT (Feng and Chen, 2024) Steps $\times$ ED&R LLM
Voicify (Vu et al., 2023) Target Page+Steps $\times$1 $\times$ ED&R Prepared dataset
UGIF (Venkatesh et al., 2022), Macro (Huang et al., 2024) Function $\times$2 $\times$ ED&R Prepared dataset
AXNav (Taeb et al., 2024) Function ED&R LLM
AutoDroid (Wen et al., 2024) Function ED&R Prepared UI graph
GPTVoiceTasker (Vu et al., 2024) Function UI Expanding UI Graph
RTA (Zhang et al., 2023b) Task-related texts ED&R LLM
CogAgent (Hong et al., 2023a) Task-related texts ED&R Fine-tuned LLM
VisionTasker (Song et al., 2024) Task-related texts ED&R LLM, prepared dataset
Mobile-Agent (Wang et al., 2024) Task-related texts ED&R LLM
Prompt2Task Task-related texts UI Online tutorials,
accumulated knowledge
  • ED&R: Extend the dataset and retrain models to handle errors.
  • UI: User intervention.
  • 1 Only pages with deep links are supported.
  • 2 Only functions in the dataset are supported.
2. Output: The output environment for task automation is dynamic, which means that the operation sequence ($o_{1:n}$) depends not only on the textual instructions ($i_{1:n}$) but also on the operating environment. Factors like smartphone model (Al-Razgan et al., 2012), app version (Venkatesh et al., 2022), and page state (Huang et al., 2023) can all influence the operation sequence. For example, as shown in Figure 3, when a user wants to post a moment, varying page states may require different operations.
3. Error Handling: Successfully executing a multi-step automation task requires that each step be error-free, which is a strict condition. Besides the two challenges mentioned above, many other factors can lead to errors, such as missing accessibility labels on GUIs (Xie et al., 2020; White et al., 2019; Wang et al., 2023), inadequate semantic recognition of GUI elements (Dixon et al., 2014; Passini et al., 2008; Liu et al., 2018; Lee et al., 2022), and missing parameters during execution (Ni et al., 2020). These issues can result in unpredictable outcomes and significantly compromise the reliability of the task automation system. While we acknowledge that data-driven models have the potential to improve accuracy with sufficient training data in the future, we are more eager to optimize the model’s performance in a controlled and targeted manner for current usage scenarios. Usually, a simple and minor correction is sufficient to ensure the success of a task (Li et al., 2024a).
4. Knowledge: Since the process requires the infusion of knowledge, ensuring reliable knowledge acquisition is critical. While LLMs are promising for planning and reasoning, they often lack the necessary domain-specific knowledge, leading to errors or hallucinations (Ji et al., 2023), especially with uncommon or deeply hidden functions in applications (Vu et al., 2024). Additionally, LLMs are often inefficient (Wang et al., 2023), making the automation process even slower than manual execution, which users find frustrating (Zhao et al., 2015). Recent works have started to recognize the importance of domain-specific knowledge, such as continuously constructing UI graphs to enhance task planning (Vu et al., 2024). We believe a more systematic and comprehensive application of domain-specific knowledge is needed—one that ensures the knowledge required for task automation is not only accurate but also efficiently extracted and applied, improving both reliability and speed throughout the entire process.
Considering the aforementioned challenges, prior work has proposed various solutions, each tackling some aspects of these issues to varying degrees. As shown in Table 1, we identify significant room for improvement in their practical effectiveness when applied in real-world scenarios. To this end, we have designed and implemented Prompt2Task, which provides targeted solutions to the challenges outlined above.
4. Prompt2Task
To address the above challenges and align with the components of the Prompt2Task model, we propose Prompt2Task, as shown in Figure 4. We have divided the process into three stages:
@@ 4. Prompt2Task
Figure 4. System design of Prompt2Task. The blue solid lines represent the main flow of the task automation process, the grey dashed lines represent optional user intervention, the purple lines represent knowledge accumulation after task completion and its optimization to each stage, and the red solid line denotes instances where the agent seeks user assistance. Each stage’s workflow details are shown in Figure 5, 7, 8.
Information Collection: To tackle the challenge posed by inputs, Prompt2Task extracts useful information and underlying intentions from the textual prompt $P$. It supplements missing information in the user’s expression through predictive analysis, asking the user or retrieving external knowledge. This stage prepares all relevant natural language information: the function description $F$ and step description $s_{1:n}$.
Instruction Generation: Employing the collected information, this stage strategizes the task’s progression. Given the instruction definition, Prompt2Task converts the step description into a list of standardized instructions $i_{1:n}$.
Operation Mapping: To address challenges from the dynamic output environment, such as rapidly updated pages, Prompt2Task should make adaptive decisions beyond pre-defined instructions. Specifically, it includes a deep understanding of the real-time environment, flexible grounding, and assessment to enhance the reliability of operations. This stage finalizes the generation of the operation sequence $o_{1:n}$.
To optimize the system in a controlled and targeted manner, Prompt2Task is designed to seek user assistance when errors or uncertainties arise, ensuring the task remains on track. The trade-off here is balancing user oversight to enhance system effectiveness while minimizing user burden. Prompt2Task achieves this by making the results of each stage transparent and requesting user help only when necessary. Even if users are not willing to intervene, Prompt2Task can autonomously complete the entire process.
In terms of knowledge, we incorporate online tutorials to expand the task range as open-ended as possible and enhance the reliability of planning. The accumulated knowledge from past tasks, including context and execution information, is processed and stored to improve Prompt2Task’s future reasoning capabilities and efficiency.
Throughout the above three stages, different types of intelligence and knowledge are applied to achieve the corresponding goals. Therefore, we utilize a multi-agent system framework. Each agent, based on their assigned sub-tasks, is responsible for specific knowledge retrieval, tool utilization, decision-making, and action execution. These agents collaborate to complete the overall task. Figure 4 illustrates how Prompt2Task is partitioned into specialized agents and a unified knowledge base. We will delve into each in the following subsections.
@@ 4.1. Multi-Agent Design
Drawing inspiration from the prior agent system frameworks (Panait and Luke, 2005; Dorri et al., 2018; Talebirad and Nadiri, 2023; Hong et al., 2023b), each agent within Prompt2Task is composed of several key components: memory, tools, strategy, and actions, which support the agent’s fundamental functions. Additionally, each agent is equipped with a user interface that allows users to observe, modify, or augment agent decisions, as shown in Figure 4. The interface is not just for monitoring but aims to foster human-agent collaboration.
We have designed five distinct agents specifically tailored to the three stages of the process.
  • Information Collection (Figure 5):
    • Analysis Agent: Extracts and analyzes information implicit in the textual prompt to construct a complete function description $F$.
    • Retrieval Agent: Acquires external knowledge relevant to the function description and builds an enriched step description $s_{1:n}$.
  • Instruction Generation (Figure 7):
    • Parsing Agent: Transforms the collected information into a list of instructions $i_{1:n}$.
  • Operation Mapping (Figure 8):
    • Grounding Agent: Predicts operations $o_{1:n}$ on the smartphone based on the generated instructions.
    • Assessment Agent: Reviews the predicted operations, evaluates if user intervention is required, and eventually executes operations on smartphones.
@@
Figure 5. Information collection workflow. Task-related functions and step descriptions are collected by the Analysis Agent and Retrieval Agent. The colors of the components within the prompt indicate their sources, which are the same color elements pointing to the prompt with the same color arrows. Gray arrows represent the normal task flow, while purple arrows indicate the accumulation of knowledge.
@@ 4.1.1. Analysis Agent
The agent aims to analyze and confirm task-related information from the textual prompts.
The prompts may include descriptions of the function or specific steps in the model, as outlined in Section 3.1. The function description must be confirmed and refined from the input. The original function descriptions are often inforaml and may lack essential context, such as specific application names or phone models. Figure 6 demonstrates how the agent interacts with users to refine these descriptions. Utilizing the LLM, the agent successively adopts three key strategies:
@@
  • Normalization: Users often describe the desired outcome without accurately naming the specific function. To address this, the agent rephrases the original input to reduce colloquial and inaccurate expressions.
  • Automated Context Enrichment: The agent utilizes the LLM to enhance the user’s original input by incorporating the user’s record from the context library (detailed in Section 4.2.2). The record is directly added in the prompt to the LLM, enabling the generation of several potential candidates for the user to choose from.
  • Interactive Context Enrichment: If the agent determines that relevant context information is still missing after the automated enrichment, it queries the user to obtain the missing details. The agent continues to chat with the user until it determines that the function description is sufficiently clear.
After refining and enriching the function description, the user confirms it via chatting (Figure 11). The agent then bundles the function description, along with step descriptions (if available), and forwards them to the Retrieval Agent for subsequent processing.
@@ 4.1.2. Retrieval Agent
The Analysis Agent can ensure a clear function description, but it may fall short when it comes to providing detailed step descriptions, which the prompt might not always cover. Addressing this void, the Retrieval Agent is responsible for preparing a complete step description. For efficiency, the agent prioritizes the following strategies:
@@
  • Otherwise, the agent searches the historical task repository (detailed in Section 4.2.1) for matching the most relevant historical task model, using function description as the keyword. If a match is found, the agent will use the matched task model, which has been previously used within our system and is deemed reliable.
If the user is dissatisfied or no match is found, the agent searches for relevant tutorials on the Internet. The agent uses search engines, with the function description as a query. The tutorial texts from the fetched web pages are refined using SimpRead [^2] https://simpread.pro/’s rule library and summarized by the LLM. As illustrated in Equation 1, the score for each tutorial is calculated based on criteria that include their original search ranking, the data source, and textual quality assessed by the LLM (Chen et al., 2023). Notably, the alignment with the original step description analyzed from the user’s prompt is also a factor to consider, which is also assessed by the LLM. The constant parameters are empirically determined. Tutorials are ranked according to their score from highest to lowest.
(1) $Score(Tutorial_{i})$ $=c_{0}\times originalRank_{i}$
$+c_{1}\times weight(source_{i})$
$+c_{2}\times textQuality_{i}$
$+c_{3}\times matched(i,oriStepDescription)$
As shown in Figure 11, Prompt2Task presents the top 5 highest-ranking tutorials to users, automatically selecting the highest-ranked tutorial in cases where users do not make a choice. By this stage, both the function and step description for the task automation are prepared in place.
Figure 7. Instruction generation workflow. Task-related instructions are generated by the Parsing Agent. The colors of the components within the prompt indicate their sources, which are the same color elements pointing to the prompt with the same color arrows. Gray arrows represent the normal task flow, while purple arrows indicate the accumulation of knowledge.